OpenStudy (anonymous):

How do you work these types of problems out: Problem type A: √100 Problem type B: -√16 Problem type C: 1.1 ^ 2 Problem type D: (-18) ^2 Problem type E: √36 + √64

OpenStudy (anonymous):

\[100=10\times 10\]

OpenStudy (anonymous):

okay.

OpenStudy (anonymous):

\[10\times10=10^2\]

OpenStudy (anonymous):

alright.

OpenStudy (anonymous):

type b?

OpenStudy (anonymous):

@Jonask

OpenStudy (anonymous):

please?

OpenStudy (aroub):

For B: \[4^2=16\] So th answer is -4

OpenStudy (anonymous):

\[-(\pm4)=\mp4\]

OpenStudy (anonymous):

i really dont get it

OpenStudy (aroub):

For C: It's easy.. You just do 1.1 * 1.1

OpenStudy (anonymous):

i do not get b.

OpenStudy (aroub):

Aoh, You don't know the concept of square roots right?

OpenStudy (anonymous):

no.

OpenStudy (anonymous):

\[\sqrt{16}= \pm4\] \[-\sqrt{16}= -(\pm)4 = \mp4\]

OpenStudy (anonymous):

B. is basically saying -1 * square root 16. Think what number multiplied by itself equals 16. Then multiply it by -1.

OpenStudy (anonymous):

can i give yo the whole list of problems? this is really difficult for me.

OpenStudy (anonymous):

a^2 = a*a a^2 = b The square root of "b" would be "a". "a" was squared in order to equal "b" so it's the sqaure root.

OpenStudy (anonymous):

\[\sqrt{16} = \pm4\] but in general only principle value (i.e. +4) is taken. so if you consider only principle value then it is \[-(\sqrt{16})=-(4)=-4\]

OpenStudy (aroub):

Okay lets take B as an example. Find \[-\sqrt{16}\] Square of 16= 4^2=16 \[-\sqrt[2]{4^2}\] The square root always has the index 2 so just imagine that you're cancelling the index by the 2. 

OpenStudy (aroub):

Do you get it?!

OpenStudy (anonymous):

-sigh- no.

OpenStudy (anonymous):

couldnt you do a simpler explanation?

OpenStudy (aroub):

Simpler than that? What don't you get?

OpenStudy (anonymous):

what grade are you in,aroub?

OpenStudy (anonymous):

When you square a number (ex 10^2) you multiply it by itself: 10 * 10 10 * 10 = 100 so 10^2 = 100 The square root of 100 would be 10 because it was squared to equal 100. Do you understand? Problem type A: √100 10 * 10 = 100 so 10 is the square root. Problem type B: -√16 -1 * √16 4 * 4 = 16 so 4 is the suare root. 4 * -1 = -4 Problem type C: 1.1 ^ 2 What is 1.1 * 1.1? Problem type D: (-18) ^2 What is -18 * -18? Problem type E: √36 + √64 What # times itself equals 36? What # times itself equals 64? Add them together.

OpenStudy (anonymous):

root = begining/source

OpenStudy (anonymous):

i will try.

OpenStudy (anonymous):

√67

OpenStudy (anonymous):

What problem were you working on and how did you get the √67 ?

OpenStudy (anonymous):

that is another problem

OpenStudy (anonymous):

i will try jazy, then message u.

OpenStudy (anonymous):

okay?

OpenStudy (anonymous):

Sounds good, although I'm not very sure on how to work out √67

OpenStudy (anonymous):

okay.